solve w/21-w/35=1/21
Answer by Adrish(44)

Solutions: Multiply the whole equation by 21, we get
w/21x 21 -w/35 x 21 =1/21 x21
w -3w/5 =1
(5w-3w)/5 =1
2w/5 =1
By cross multiplying we get
2w= 5
w=5/2